# Glyph-ByT5: A Customized Text Encoder for Accurate Visual Text Rendering

### v1

@@ Abstract
[Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
@@ 3.1 Glyph-ByT5: Customized Glyph-Aligned Character-Aware Text Encoder for Design-text Generation
Figure 2: Illustrating the scheme of glyph augmentation. (a) original glyph. (b) character replacement (Happy $\to$ Hdppy). (c) character repeat (Happy $\to$ Happpppy). (d) character drop (Happy $\to$ Hapy). (e) character add (Graduation $\to$ Gradumation). (f) word replacement (Graduation $\to$ Gauatikn). (g) word repeat (Kim $\to$ Kim Kim). (h) word drop (Happy Graduation Kim $\to$ Graduation).
Refer to caption (a) Refer to caption (b) Refer to caption (c) Refer to caption (d) Refer to caption (e) Refer to caption (f) Refer to caption (g) Refer to caption (h)
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Figure 2: Illustrating the scheme of glyph augmentation. (a) original glyph. (b) character replacement (Happy $\to$ Hdppy). (c) character repeat (Happy $\to$ Happpppy). (d) character drop (Happy $\to$ Hapy). (e) character add (Graduation $\to$ Gradumation). (f) word replacement (Graduation $\to$ Gauatikn). (g) word repeat (Kim $\to$ Kim Kim). (h) word drop (Happy Graduation Kim $\to$ Graduation).
@@
Creating Scalable and Accurate Glyph-Text Dataset To enable the training of the customized glyph-aware text encoder, we first create a high-quality glyph-text dataset, denoted as $\mathcal{D}$, consisting of approximately $\sim 1$ million pairs of synthetic data $\{\mathsf{I}_{\rm{glyph}},\mathsf{T}_{\rm{text}}\}$. This dataset was developed with the improved graphic render introduced in the recent work by [13]. We construct the initial glyph image set based on the original typographic attributes (including font types, colors, sizes, positions, and others) found in the crawled graphic design images. We compile a large text corpus that can be used to enrich the glyph image set by replacing the words with random text sampled from the corpus. Additionally, we randomly modify the font types and colors within each text box to further enlarge the dataset. Our glyph-text dataset $\mathcal{D}$ encompasses nearly $\sim 512$ different font types and $\sim 100$ distinct font colors. To ensure the glyph-aligned text encoder focuses on only the difference on the visual text, we all use black colored background by default.
We present the example of glyph prompts corresponding to the glyph image shown in Figure 3 (a), detailing font types, colors, and text, as illustrated follows: {Text “oh! the places you’ll go!” in [font-color-39], [font-type-90]. Text “Happy Graduation Kim!” in [font-color-19] [font-type-181]}. In this process, special tokens are utilized to denote font colors and types. Prior to inputting it into the Glyph-ByT5 text encoder, we preprocess the prompt text by substituting special tokens, like the token ‘[font-color-39]’, with a series of global embeddings from the enriched codebook. We have conducted experiments on the Glyph-Text datasets at three distinct scales, expanding from 100K to 500K, and up to 1M. In the future, we aim to significantly expand our datasets, scaling up to 100M given access to more computing resources.
@@ 3.2 Glyph-SDXL: Augmenting SDXL with Glyph-ByT5 for Design Image Generation
Refer to caption (a) Glyph-Alignment Pre-training Refer to caption (b) Region-wise Multi-Text-Encoder Fusion
@@ 4.2 VisualParagraphy Benchmark
Figure 6: The user study results. Higher score means better results.
@@
Table 2: Comparison on SimpleBench, CreativeBench, and MARIO-Eval.
Table 3: Effect of using different pre-trained visual encoder.
Table 5: Effect of Glyph Augmentation Ratio.
Table 6: Effect of the ByT5-to-SDXL mapper within Glyph-SDXL.
@@
Table 2: Comparison on SimpleBench, CreativeBench, and MARIO-Eval.
Table 3: Effect of using different pre-trained visual encoder.
Table 5: Effect of Glyph Augmentation Ratio.
Table 6: Effect of the ByT5-to-SDXL mapper within Glyph-SDXL.
@@ 4.6 Ablation Experiments
Mapper, Scaling Glyph-Text Dataset and Text Encoder Size, and More Table 8 shows the importance of using the ByT5-to-SDXL mapper to align the gap. Table 8 and Table 8 verify the benefits of scaling up the glyph-text dataset size and text encoder size. We provide more ablation experiments of the Glyph-SDXL-Scene in the supplementary material.
@@ A. Dataset Statistics
Refer to caption (a) Refer to caption (b) Refer to caption (c) Refer to caption (d) Refer to caption (e) Refer to caption (f)
@@ B. Training Settings
Table 10: Glyph-SDXL model training hyper-parameters.
@@
Table 10: Glyph-SDXL model training hyper-parameters.
@@ C. Typography Editing with Region-wise SDEdit
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Figure 9: Illustrating the Region-wise SDEdit pipeline.
@@
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption (a) $t_{0}=800$ Refer to caption (b) $t_{0}=700$ Refer to caption (c) $t_{0}=600$ Refer to caption (d) $t_{0}=500$ Refer to caption (e) $t_{0}=400$ Refer to caption (f) original
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption (a) $t_{1}=600$ Refer to caption (b) $t_{1}=500$ Refer to caption (c) $t_{1}=400$ Refer to caption (d) $t_{1}=300$ Refer to caption (e) $t_{1}=200$ Refer to caption (f) $t_{1}=100$
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Figure 11: Effect of different choices of $t_{0}$ for Region-wise SDEdit.
Figure 12: Effect of different choices of $t_{1}$ for Region-wise SDEdit.
@@
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
@@
E. Ablation on other Text Encoder Fusion Scheme
One might question the effectiveness of a basic approach that concatenates the text embeddings from different text encoders. The comparison results are detailed in Table 11. Empirically, we find that this baseline underperforms significantly due to the substantial discrepancies among the text encoders.
@@ G. Detailed Prompt List
Fig 1, Row 2, Col1 Background: Instagram Posts. The image features a stack of pancakes with syrup and strawberries on top. The pancakes are arranged in a visually appealing manner, with some pancakes placed on top of each other. The syrup is drizzled generously over the pancakes, and the strawberries are scattered around, adding a touch of color and freshness to the scene. The overall presentation of the pancakes is appetizing and inviting. Tags: brown, peach, grey, modern, minimalist, simple, colorful, illustration, Instagram post, instagram, post, national pancake day, international pancake day, happy pancake day, pancake day, pancake, sweet, cake, discount, sale. Text: Text ”Get 75% Discount for your first order” in $<$color-3$>$, $<$font-25$>$. Text ”Order Now” in $<$color-0$>$, $<$font-97$>$. Text ”National Pancake Day” in $<$color-4$>$, $<$font-97$>$.
@@
Fig 1, Row 2, Col2 Background: Cards and invitations. The image features a large gray elephant sitting in a field of flowers, holding a smaller elephant in its arms. The scene is quite serene and picturesque, with the two elephants being the main focus of the image. The field is filled with various flowers, creating a beautiful and vibrant backdrop for the elephants. Tags: Light green, orange, Illustration, watercolor, playful, Baby shower invitation, baby boy shower invitation, baby boy, welcoming baby boy, koala baby shower invitation, baby shower invitation for baby shower, baby boy invitation, background, playful baby shower card, baby shower, card, newborn, born, Baby Shirt Baby Shower Invitation. Text: Text ”RSVP to +123-456-7890” in $<$color-18$>$, $<$font-100$>$. Text ”Olivia Wilson” in $<$color-99$>$, $<$font-210$>$. Text ”Baby Shower” in $<$color-53$>$, $<$font-210$>$. Text ”Please Join Us For a” in $<$color-18$>$, $<$font-100$>$. Text ”In Honoring” in $<$color-18$>$, $<$font-100$>$. Text ”23 November, 2021 — 03:00 PM Fauget Hotels” in $<$color-18$>$, $<$font-100$>$.
@@
Fig 1, Row 2, Col3 Background: Flyers. The image features a purple background with a witch flying on a broomstick, surrounded by several pumpkins. The pumpkins are scattered throughout the scene, with some positioned closer to the witch and others further away. The combination of the purple background, the witch, and the pumpkins creates a festive and spooky atmosphere. Tags: purple, orange, colorful, illustration, creative, fun, dark, bold, playful, cute, cartoon, flyer, halloween, trick or treat, costume, party, spooky, pumpkin, trick, event. Text: Text ”Games” in $<$color-27$>$, $<$font-197$>$. Text ”Costume Party” in $<$color-27$>$, $<$font-197$>$. Text ”Candies” in $<$color-27$>$, $<$font-197$>$. Text ”October 31st 6 p.m. - 9 p.m. 123 Anywhere St., Any City, ST 12345” in $<$color-51$>$, $<$font-197$>$. Text ”Treat” in $<$color-51$>$, $<$font-371$>$. Text ”or” in $<$color-51$>$, $<$font-371$>$. Text ”Trick” in $<$color-51$>$, $<$font-371$>$.
@@
Fig 1, Row 2, Col4 Background: Instagram Posts. The image features a purple witch’s hat on a pumpkin, which is placed in front of a graveyard. The pumpkin is positioned in the center of the scene, and the hat is slightly tilted to the left. There are three ghosts in the background, with one on the left side, one on the right side, and another one in the middle. The ghosts are positioned at different heights, with the one on the left being the tallest, the one in the middle being the shortest, and the one on the right being slightly taller than the middle ghost. Tags: purple, orange, yellow, illustration, halloween, halloween day, halloween party, happy halloween, pumpkins, trick or treats, spooky, haunted, event, party, festive, witch, monster, scary, ghost, instagram post. Text: Text ”Big deals” in $<$color-14$>$. Text ”31 October 2022” in $<$color-14$>$. Text ”HALLOWEEN SALE” in $<$color-57$>$, $<$font-252$>$. Text ”ONLY FOR TODAY” in $<$color-57$>$, $<$font-252$>$. Text ”50% OFF” in $<$color-57$>$, $<$font-252$>$.
@@
Fig 1, Row 1, Col4 Background: A photo of a cute squirrel holding a sign, 4k, dslr. Text: Text ”Glyph-ByT5: A Customized Text Encoder for Accurate Visual Text Rendering”.
@@
Fig 1, Row 1, Col4 Background: A man standing in the midst of a vibrant sunflower field with a mountain range in the background under a blue sky, holding a sign that reads ”Glyph-ByT5: A Customized Text Encoder for Accurate Visual Text Rendering” Vincent van Gogh style. Text: Text ”Glyph-ByT5: A Customized Text Encoder for Accurate Visual Text Rendering”.
@@
Fig 1, Row 1, Col4 Background: An intriguing scene of a blank sign standing amidst a rocky landscape, with a backdrop of a clear sky and a palm tree. Text: Text ”Words on sign are”. Text ”CORRECT”. Text ”exactly with Glyph ByT5”.
@@
Fig 1, Row 1, Col4 Background: The image shows a sign with a stylized design, featuring a bird and branches. The sign is hanging from a ceiling, and it appears to be located outside a building. The design is simple and modern, with a limited color palette that includes shades of brown and white. The bird and branches are depicted in a minimalist style, with clean lines and a lack of detail that gives the sign a contemporary feel. The sign is likely intended to provide information or direction to passersby, but the specific content of the sign is not visible in the image. Text ”Glyph ByT5: A Customized Text Encoder for Accurate”. Text ”Visual Text Rendering”.
@@
Fig 5, Row 1, Col2 Background: The image features a white background with a few plants and flowers scattered across it. There are three main plants in the scene, with one located on the left side, another in the middle, and the third on the right side. Additionally, there are two smaller plants in the upper part of the image. The plants are of various sizes and shapes, adding a sense of diversity to the scene. Text: Text ”Give yourself the same amount of care you selflessly give to others.” in $<$color-7$>$, $<$font-196$>$.
Fig 5, Row 1, Col3 Background: Instagram Posts. The image features a woman sitting in a lotus position, also known as a yoga pose, with her legs crossed and her hands resting on her knees. She is surrounded by a serene environment, with trees in the background and a sun in the sky above her. The woman appears to be meditating or practicing yoga in a peaceful outdoor setting. Tags: WHITE, BROWN, BLUE, MODERN, meditation, exercise, fitness, yoga day, poster, health, illustration, international, position, concept, relaxation, yoga, woman. Text: Text ”” letś move to find healthy”” in $<$color-2$>$, $<$font-30$>$. Text ”YOGA DAY” in $<$color-2$>$, $<$font-30$>$. Text ”International” in $<$color-2$>$, $<$font-500$>$.
Fig 5, Row 1, Col4 Background: The image features a white background with a circular frame made of colored pencils. The frame is filled with a variety of colored pencils, creating a visually appealing and artistic design. The pencils are arranged in different positions, with some overlapping and others standing alone. The combination of colors and the circular shape of the frame make the image a unique and creative piece of art. Text: Text ”I TOTALLY REMEMBERED YOUR BIRTHDAY!” in $<$color-14$>$, $<$font-101$>$.
Fig 5, Row 1, Col5 Background: Facebook Post. The image features a man wearing a black shirt and blue overalls, holding a brown box. He is standing in front of a blue background, which has a few speech bubbles scattered around. The man appears to be smiling, possibly indicating that he is happy or excited about the box he is holding. Tags: Violet, purple, illustration, illustrated, corporate, professional, Courier, delivery, parcel, package, fast, free, express, shipping, vehicle, transportation, pickup, centre, man, character. Text: Text ”EXPRESS PARCEL SHIPPING” in $<$color-0$>$, $<$font-4$>$. Text ”www.reallygreatsite.comïn $<$color-0$>$, $<$font-4$>$. Text ”tel.: +123-456-7890” in $<$color-0$>$, $<$font-4$>$. Text ”over 1000 Delivery Centres” in $<$color-0$>$, $<$font-4$>$. Text ”Online Tracking” in $<$color-0$>$, $<$font-4$>$. Text ”10% off for New Clients” in $<$color-0$>$, $<$font-4$>$. Text ”FOR BUSINESS AND INDIVIDUALS” in $<$color-123$>$, $<$font-4$>$. Text ”DELIVERY” in $<$color-123$>$, $<$font-54$>$. Text ”COURIER” in $<$color-0$>$, $<$font-54$>$.

### v2

@@ 3.1 Glyph-ByT5: Customized Glyph-Aligned Character-Aware Text Encoder for Design-text Generation
Figure 2: Illustrating the scheme of glyph augmentation. (a) original glyph. (b) character replacement (Happy $\to$ Hdppy). (c) character repeat (Happy $\to$ Happpppy). (d) character drop (Happy $\to$ Hapy). (e) character add (Graduation $\to$ Gradumation). (f) word replacement (Graduation $\to$ Gauatikn). (g) word repeat (Happy $\to$ Happy Happy). (h) word drop (Happy Graduation Amber $\to$ Graduation).
@@
Creating Scalable and Accurate Glyph-Text Dataset To enable the training of the customized glyph-aware text encoder, we first create a high-quality glyph-text dataset, denoted as $\mathcal{D}$, consisting of approximately $\sim 1$ million pairs of synthetic data $\{\mathsf{I}_{\rm{glyph}},\mathsf{T}_{\rm{text}}\}$. This dataset was developed with the improved graphic render introduced in the recent work by [13]. We construct the initial glyph image set based on the original typographic attributes (including font types, colors, sizes, positions, and others) found in the crawled graphic design images. We compile a large text corpus that can be used to enrich the glyph image set by replacing the words with random text sampled from the corpus. Additionally, we randomly modify the font types and colors within each text box to further enlarge the dataset. Our glyph-text dataset $\mathcal{D}$ encompasses nearly $\sim 305$ different OFL licenced [^2] These open-source font types are licensed under the SIL Open Font License that permits commercial usage. font types and $\sim 100$ distinct font colors. To ensure the glyph-aligned text encoder focuses on only the difference on the visual text, we all use black colored background by default.
We present the example of glyph prompts corresponding to the glyph image shown in Figure 3 (a), detailing font types, colors, and text, as illustrated follows: {Text “The way you create a better future is by studying the past.” in [font-color-127], [font-type-234]. Text “Happy Graduation Amber” in [font-color-98] [font-type-231]}. In this process, special tokens are utilized to denote font colors and types. Prior to inputting it into the Glyph-ByT5 text encoder, we preprocess the prompt text by substituting special tokens, like the token ‘[font-color-127]’, with a series of global embeddings from the enriched codebook. We have conducted experiments on the Glyph-Text datasets at three distinct scales, expanding from 100K to 500K, and up to 1M. In the future, we aim to significantly expand our datasets, scaling up to 100M given access to more computing resources.
@@ 3.2 Glyph-SDXL: Augmenting SDXL with Glyph-ByT5 for Design Image Generation
(a) Glyph-Alignment Pre-training
(b) Region-wise Multi-Text-Encoder Fusion
@@ 4.2 VisualParagraphy Benchmark
[…]
@@
Table 2: Comparison on SimpleBench, CreativeBench, and MARIO-Eval.
@@
$\leq$20 chars $\leq$20-50 chars $\leq$50-100 chars $\geq$100 chars
DINOv2 ViT-B/14 + reg $\bf{84.54}$ $\bf{84.56}$ $\bf{79.89}$ $\bf{73.29}$
CLIP ViT-B/16 $77.17$ $74.78$ $74.94$ $66.34$
ViTSTR $79.29$ $78.2$ $75.35$ $68.49$
CLIP4STR ViT-B/16 $80.38$ $79.12$ $77.08$ $69.24$
Table 3: Effect of using different pre-trained visual encoder.
@@
Table 6: Effect of the ByT5-to-SDXL mapper within Glyph-SDXL.
@@ 4.6 Ablation Experiments
Mapper, Scaling Glyph-Text Dataset and Text Encoder Size, and More Table 8 shows the importance of using the ByT5-to-SDXL mapper to align the gap. Table 8 and Table 8 verify the benefits of scaling up the glyph-text dataset size and text encoder size. We provide more ablations, as well as experiments of Glyph-SDXL-Scene in the appendix.
@@ B. Training Settings
Table 10: Glyph-SDXL model training hyper-parameters.
[…]
$\leq$20 chars $\leq$20-50 chars $\leq$50-100 chars $\geq$100 chars
GlyphControl-SDXL $65.17$ $61.04$ $45.31$ $32.75$
Glyph-SDXL $\bf{84.54}$ $\bf{84.56}$ $\bf{79.89}$ $\bf{73.29}$
Table 11: Comparison with ControlNet style model.
[…]
$\leq$20 chars $\leq$20-50 chars $\leq$50-100 chars $\geq$100 chars
SDXL-VAE $70.02$ $70.57$ $66.34$ $56.28$
DINOv2 $\bf{84.54}$ $\bf{84.56}$ $\bf{79.89}$ $\bf{73.29}$
Table 12: Comparison with aligning to the latent space of SDXL.
@@ D. Ablation on the Design-to-Scene Alignment
E. More ablation experiments
Comparison with ControlNet-style Model To validate the effectiveness of our customized text encoder against other forms of feature guidance, we fine-tune a representative ControlNet-style SDXL model, i.e., GlyphControl [30], and report results in Table 13. We can see that our Glyph-SDXL significantly performs better, and the performance gap increases with a larger number of characters. This indicates that the ControlNet-style model, conditioned on rendered text masks, suffers from fundamental limitations when handling paragraph-level dense visual text.
Comparison with Aligning Directly with SDXL Latent Space We compare our approach with directly aligning with SDXL’s latent space, i.e. using SDXL-VAE as the visual encoder. As indicated in Table 13, SDXL-VAE is weak at extracting reliable visual text features and lags behind DINOv2 significantly.
Ablation on other Text Encoder Fusion Scheme One might question the effectiveness of a basic approach that concatenates the text embeddings from different text encoders. The comparison results are detailed in Table 13. Empirically, we find that this baseline underperforms significantly due to the substantial discrepancies among the text encoders.
@@ G. Detailed Prompt List
Fig 1, Row 2, Col1 Background: Instagram Posts. The image features a stack of pancakes with syrup and strawberries on top. The pancakes are arranged in a visually appealing manner, with some pancakes placed on top of each other. The syrup is drizzled generously over the pancakes, and the strawberries are scattered around, adding a touch of color and freshness to the scene. The overall presentation of the pancakes is appetizing and inviting. Tags: brown, peach, grey, modern, minimalist, simple, colorful, illustration, Instagram post, instagram, post, national pancake day, international pancake day, happy pancake day, pancake day, pancake, sweet, cake, discount, sale. Text: Text ”Get 50% Discount for your first order” in $<$color-3$>$, $<$font-97$>$. Text ”Have a try!” in $<$color-0$>$, $<$font-97$>$. Text ”Celebrate with Pancakes!” in $<$color-4$>$, $<$font-97$>$.
@@
Fig 1, Row 2, Col2 Background: Cards and invitations. The image features a large gray elephant sitting in a field of flowers, holding a smaller elephant in its arms. The scene is quite serene and picturesque, with the two elephants being the main focus of the image. The field is filled with various flowers, creating a beautiful and vibrant backdrop for the elephants. Tags: Light green, orange, Illustration, watercolor, playful, Baby shower invitation, baby boy shower invitation, baby boy, welcoming baby boy, koala baby shower invitation, baby shower invitation for baby shower, baby boy invitation, background, playful baby shower card, baby shower, card, newborn, born, Baby Shirt Baby Shower Invitation. Text: Text ”RSVP to +123-456-7890” in $<$color-18$>$, $<$font-100$>$. Text ”John Doe” in $<$color-99$>$, $<$font-210$>$. Text ”Baby Shower” in $<$color-53$>$, $<$font-210$>$. Text ”Please Join Us For a” in $<$color-18$>$, $<$font-100$>$. Text ”In Honoring” in $<$color-18$>$, $<$font-100$>$. Text ”01 July, 2024 — 00:00 PM Grand Central Hotel” in $<$color-18$>$, $<$font-100$>$.
@@
Fig 1, Row 2, Col3 Background: Flyers. The image features a purple background with a witch flying on a broomstick, surrounded by several pumpkins. The pumpkins are scattered throughout the scene, with some positioned closer to the witch and others further away. The combination of the purple background, the witch, and the pumpkins creates a festive and spooky atmosphere. Tags: purple, orange, colorful, illustration, creative, fun, dark, bold, playful, cute, cartoon, flyer, halloween, trick or treat, costume, party, spooky, pumpkin, trick, event. Text: Text ”Games” in $<$color-27$>$, $<$font-197$>$. Text ”Costume Party” in $<$color-27$>$, $<$font-197$>$. Text ”Candies” in $<$color-27$>$, $<$font-197$>$. Text ”Warmly welcome to join us at our new playground on October 31st from 6 p.m. to 9 p.m.” in $<$color-51$>$, $<$font-197$>$. Text ”Treat” in $<$color-51$>$, $<$font-371$>$. Text ”or” in $<$color-51$>$, $<$font-371$>$. Text ”Trick” in $<$color-51$>$, $<$font-371$>$.
@@
Fig 1, Row 2, Col4 Background: Instagram Posts. The image features a purple witch’s hat on a pumpkin, which is placed in front of a graveyard. The pumpkin is positioned in the center of the scene, and the hat is slightly tilted to the left. There are three ghosts in the background, with one on the left side, one on the right side, and another one in the middle. The ghosts are positioned at different heights, with the one on the left being the tallest, the one in the middle being the shortest, and the one on the right being slightly taller than the middle ghost. Tags: purple, orange, yellow, illustration, halloween, halloween day, halloween party, happy halloween, pumpkins, trick or treats, spooky, haunted, event, party, festive, witch, monster, scary, ghost, instagram post. Text: Text ”Big promotion” in $<$color-14$>$. Text ”31 October 2024” in $<$color-14$>$. Text ”HALLOWEEN SALE” in $<$color-57$>$, $<$font-252$>$. Text ”ONCE IN A YEAR” in $<$color-57$>$, $<$font-252$>$. Text ”60% OFF” in $<$color-57$>$, $<$font-252$>$.
@@
Fig 1, Row 3, Col1 Background: A photo of a cute squirrel holding a sign, 4k, dslr. Text: Text ”Glyph-ByT5: A Customized Text Encoder for Accurate Visual Text Rendering”.
@@
Fig 1, Row 3, Col2 Background: A man standing in the midst of a vibrant sunflower field with a mountain range in the background under a blue sky, holding a sign that reads ”Glyph-ByT5: A Customized Text Encoder for Accurate Visual Text Rendering” Vincent van Gogh style. Text: Text ”Glyph-ByT5: A Customized Text Encoder for Accurate Visual Text Rendering”.
@@
Fig 1, Row 3, Col3 Background: An intriguing scene of a blank sign standing amidst a rocky landscape, with a backdrop of a clear sky and a palm tree. Text: Text ”Words on sign are”. Text ”CORRECT”. Text ”exactly with Glyph ByT5”.
@@
Fig 1, Row 3, Col4 Background: The image shows a sign with a stylized design, featuring a bird and branches. The sign is hanging from a ceiling, and it appears to be located outside a building. The design is simple and modern, with a limited color palette that includes shades of brown and white. The bird and branches are depicted in a minimalist style, with clean lines and a lack of detail that gives the sign a contemporary feel. The sign is likely intended to provide information or direction to passersby, but the specific content of the sign is not visible in the image. Text ”Glyph ByT5: A Customized Text Encoder for Accurate”. Text ”Visual Text Rendering”.
@@
Fig 5, Row 1, Col2 Background: The image features a white background with a few plants and flowers scattered across it. There are three main plants in the scene, with one located on the left side, another in the middle, and the third on the right side. Additionally, there are two smaller plants in the upper part of the image. The plants are of various sizes and shapes, adding a sense of diversity to the scene. Text: Text ”Give yourself the same amount of attention and warmth you selflessly give to others.” in $<$color-7$>$, $<$font-196$>$.
Fig 5, Row 1, Col3 Background: Instagram Posts. The image features a woman sitting in a lotus position, also known as a yoga pose, with her legs crossed and her hands resting on her knees. She is surrounded by a serene environment, with trees in the background and a sun in the sky above her. The woman appears to be meditating or practicing yoga in a peaceful outdoor setting. Tags: WHITE, BROWN, BLUE, MODERN, meditation, exercise, fitness, yoga day, poster, health, illustration, international, position, concept, relaxation, yoga, woman. Text: Text ”” let’s get moving for a healthy body!”” in $<$color-2$>$, $<$font-30$>$. Text ”BEAUTY OF YOGA” in $<$color-2$>$, $<$font-30$>$. Text ”Experience” in $<$color-2$>$, $<$font-500$>$.
Fig 5, Row 1, Col4 Background: The image features a white background with a circular frame made of colored pencils. The frame is filled with a variety of colored pencils, creating a visually appealing and artistic design. The pencils are arranged in different positions, with some overlapping and others standing alone. The combination of colors and the circular shape of the frame make the image a unique and creative piece of art. Text: Text ”I TOTALLY REMEMBERED YOUR BIRTHDAY DEAR FRIEND!” in $<$color-14$>$, $<$font-101$>$.
Fig 5, Row 1, Col5 Background: Facebook Post. The image features a man wearing a black shirt and blue overalls, holding a brown box. He is standing in front of a blue background, which has a few speech bubbles scattered around. The man appears to be smiling, possibly indicating that he is happy or excited about the box he is holding. Tags: Violet, purple, illustration, illustrated, corporate, professional, Courier, delivery, parcel, package, fast, free, express, shipping, vehicle, transportation, pickup, centre, man, character. Text: Text ”EXPRESS PARCEL SHIPPING” in $<$color-0$>$, $<$font-4$>$. Text ”awesome design website” in $<$color-0$>$, $<$font-4$>$. Text ”tel.: +123-456-7890” in $<$color-0$>$, $<$font-4$>$. Text ”over 1000 Delivery Centres” in $<$color-0$>$, $<$font-4$>$. Text ”Online Tracking” in $<$color-0$>$, $<$font-4$>$. Text ”10% off for New Clients” in $<$color-0$>$, $<$font-4$>$. Text ”FOR BUSINESS AND INDIVIDUALS” in $<$color-123$>$, $<$font-4$>$. Text ”DELIVERY” in $<$color-123$>$, $<$font-54$>$. Text ”COURIER” in $<$color-0$>$, $<$font-54$>$.